OpenStudy (anonymous):

Suppose that x = ln(A) and y = ln(B). Write the following expressions in terms of x and y ln(A-B)

OpenStudy (reemii):

log(A/B) = log(A) - log(B). but.. log(A-B)... is there really a formula for this?

OpenStudy (anonymous):

some new law i never saw before!

OpenStudy (reemii):

in the statement, did you mean ln(x-y) ?

OpenStudy (reemii):

@xlovely_Lizardx

OpenStudy (anonymous):

I already tried that it's not right Campbell_st

OpenStudy (anonymous):

you can't do it, that is why

OpenStudy (anonymous):

yeah, I'm completely stumped

OpenStudy (anonymous):

that's not an option though :0

OpenStudy (anonymous):

I didn't think I could either but my professor didn't give me the option to put no solution or what ever

OpenStudy (reemii):

\(\ln(e^x-e^y)\) that's all there is to say about this expression..

OpenStudy (reemii):

(if \(x>y\))

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

that's all I got

OpenStudy (anonymous):

and "no" "no solution" "none" or "undefined" is not aloud as a possible answer

OpenStudy (reemii):

ok then ln(e^x - e^y) is the answer, because there's no formula for ln(a-b)

OpenStudy (anonymous):

how did you find that?

OpenStudy (reemii):

from x = ln(A), y = ln(B). you already found that A=e^x and B=e^y (see the answer to the question "e)" )

OpenStudy (anonymous):

oh, that makes sense